# Pair Production in time-dependent Electric field at Finite times

### v5

@@ V Conclusion
We have conducted a detailed analysis of the electron-positron pair production from the vacuum under the influence of a time-dependent Sauter pulse electric field. We calculated the single-particle distribution function using the exact analytic solution for the mode function. We develop an analytical theory that is valid for the finite times $t>\tau.$ In particular, we find an analytical expression for the distribution function in the power series of the small parameter $(1-y).$ The interesting dynamical features of the momentum distribution function at finite times are attributed to the function that appears in this expansion. We have further elaborated on this aspect of the finite-time behavior of the longitudinal and transverse momentum spectra.
@@
We analyzed the temporal evolution of the single-particle distribution function. Our investigation reveals that the process of transition from initially virtual particles to real particles occurs in three distinct stages, which crucially depend on the longitudinal and transverse momentum components. Moreover, we have quantified the initiation of the REPP stage, determined by the momentum value, with higher momenta resulting in narrower oscillations in the transient region. We meticulously examined the LMS and TMS to understand the momentum-dependent behavior during these stages. In the LMS, one interesting feature at the beginning of the REPP stage is that the spectrum exhibits an oscillating structure as imprints of a quantum signature at the finite time when the electric field is nearly zero. The two-peaked structure in which the central Gaussian-peak structure has onset oscillation, and this quantum interference pattern evolves and fades away. Based on this observation, we identified three distinct time scales associated with this behavior. Additionally, we investigated the impact of different electric field strengths ($E_{0}$) on the duration of the interference pattern’s formation and disappearance. In the multiphoton regime, for $\gamma=2.5$, we explored the time evolution of LMS in different stages of pair production, and the spectrum shows the splitting of a smooth uni-modal structure into a multi-modal Gaussian structure near the REPP region, after which it merges into a single peak Gaussian profile at the asymptotic time limit.
@@
Utilizing the approximate expressions for the momentum distribution function, we showed that the LMS structure arises from three distinct functional behaviors. The second-order term dominantly governs the early times, while combining the first and zeroth-order terms leads to the central peak structure with onset oscillation. The first-order term contributes to oscillations, owing to the presence of a Gaussian envelope with oscillation, but its amplitude diminishes with time, resulting in a smooth spectrum profile in late time.
@@
We examined the role of vacuum polarization $u(p_{\parallel},t)$, and its counterpart $v(p_{\parallel},t)$ by plotting their time evolution. In the REPP region, both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ exhibit nearly identical oscillations with the same amplitude, and this oscillation amplitude decays for higher $p_{\parallel}$ values. This observation implies that the qualitative nature of the vacuum polarization function is significantly influenced by the longitudinal momentum. To elucidate this further, we plotted the LMS of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ These plots reveal two distinct patterns emerging at the start of the REPP stage. One pattern exhibits a Gaussian-like structure, while the other displays a deformed oscillating profile within the Gaussian envelope. These oscillations, exhibiting varying amplitudes, were confined to a small window of longitudinal momentum ($-1<p_{\parallel}<1$).Over time, the regular oscillations of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ became balanced, as explicitly observed in the plot of the LMS of the created particles for late time, where the oscillating behavior was absent due to the equilibrium between $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ As we recognize the pivotal roles of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ functions in accelerating (creation of $e^{-}$) and decelerating (annihilation of $e^{-}$) electrons, respectively, the observed oscillations in the $f(p_{\parallel},t)$ at finite times can be attributed to the characteristic behavior imprints of these functions, resulting from the acceleration and deceleration of electrons in momentum representation, which manifests as the observed oscillating structure. Next, we also discuss the influence of transverse momentum on LMS, which diminishes the value of $f(p_{\parallel})$ and smooths out the oscillation for higher $p_{\perp}$ values. Moreover, we emphasized that these oscillations, seen in the late REPP stage as transverse momentum increases and also impact the formation time of pairs.
@@
Finally, we study the dynamics of TMS, uncovering interesting results that have received limited attention from other authors, particularly in the context of the Sauter pulsed electric field. We observe fluctuating substructures that undergo regular changes in the QEPP and transient region depending on the momentum value. This behavior of substructures during this stage is also influenced by longitudinal momentum, resulting in distinct substructures in the spectrum at a given time. This finding indicates that the longitudinal and transverse momentum is related to each other during particle-antiparticle formation quantitatively and qualitatively. At the beginning of the REPP region, that substructure disappears. The shape of the spectrum becomes an invariable Gauss-like profile and is not affected by the longitudinal momentum, but its peak value still depends on it during the REPP stage. This contrast highlights the dynamic evolution of substructures in earlier stages, contrasting with the smooth Gaussian-like profile observed in the REPP stage. A similar intriguing behavior in the transverse momentum distribution has been observed in the works by Krajewska et al. Krajewska and Kamiński (2019) and Bechler et al. Bechler et al. (2023), depending on the specifics of the electric field parameters. These studies highlight a fascinating characteristic of the transverse distribution function peak that occurs at non-zero $p_{\perp}$ value depending on the frequency. This observation prompts questions about the feasibility of pair formation through tunneling theory, particularly considering the significant distribution function values at non-zero transverse momentum for low frequencies. For our case, the spectrum also shows the occurrence of a peak at non-zero transverse momentum signifies the creation of pairs through the multiphoton process at finite times.
Although studying the LMS and TMS of created quasi-particles at various dynamical stages shows a critical dependence on longitudinal and transverse momentum, it also brings attention to the intriguing physics behind the behavior of the vacuum in this finite-time scenario for the pair production process.
VI Acknowledgments
@@
Deepak gratefully acknowledge the financial support from Homi Bhabha National Institute (HBNI) for carrying out this research work.
Appendix A Appendixes
A.1 Evaluation approximate analytical expression for the distribution function
@@
$=\frac{\Gamma(c)\Gamma(c-a-b)}{\Gamma(c-a)\Gamma(c-b)}\Biggl{(}1+\frac{ab}{1+a+b-c}(1-y)+\frac{(a+1)(b+1))}{(2+a+b-c)}\frac{(1-y)^{2}}{2!}+...\Biggr{)}$
$+(1-y)^{(c-a-b)}\frac{\Gamma(c)\Gamma(a+b-c)}{\Gamma(a)\Gamma(b)}\Biggl{(}1+\frac{(c-a)(c-b)}{(1+c-a-b)}(1-y)+...\Biggr{)}$ (143)
@@
$\Gamma(\textrm{i}z)\overline{\Gamma(\textrm{i}z)}=\frac{\pi}{z\sinh{(\pi z)}},$ $|\Gamma(1+\textrm{i}z)|^{2}=\frac{\pi z}{\sinh{(\pi z)}},$ $|\Gamma(\frac{1}{2}+\textrm{i}z)|^{2}$ $=\frac{\pi}{\cosh{(\pi z)}}.$ (152)
where $\overline{\Gamma(z)}$ is complex conjugate of the $\Gamma(z).$
@@
When computing expressions like $\Gamma_{1}\overline{\Gamma_{2}}$, approximate methods prove advantageous. A frequently utilized technique entails utilizing Stirling’s formula for the Gamma function, offering a simpler yet effective approach to assess the desired expression.
@@
So,
$\Gamma_{1}\overline{\Gamma_{2}}$ $=\left(\frac{\Gamma(c)\Gamma(c-a-b-1)}{\Gamma(c-a)\Gamma(c-b)}\right)\overline{\Bigl{(}\frac{\Gamma(c)\Gamma(a+b-c)}{\Gamma(a)\Gamma(b)}}\Bigr{)}$ (156)
@@
$\Gamma_{1}\overline{\Gamma_{2}}$ $\approx|\Gamma_{1}\overline{\Gamma_{2}}|e^{\textrm{i}\mathrm{\varrho}}$ (157)
@@
$|\Gamma_{1}\overline{\Gamma_{2}}|$ $=\frac{\omega_{0}}{2\omega_{1}}\frac{\exp{(\pi\tau/2(\omega_{0}-\omega_{1}+2E_{0}\tau))}}{\sinh{(\pi\tau\omega_{0})}\sqrt{1+\omega_{1}^{2}\tau^{2}}}\sqrt{\frac{((\omega_{0}+2E_{0}\tau)^{2}-\omega_{1}^{2})}{(\omega_{1}^{2}-(2E_{0}\tau-\omega_{0})^{2})}},$ (158)
@@
We aim to express $f(\bm{p},y)$ in a series involving $(1-y)$. We can then consider truncating higher-order terms to simplify the analysis while still capturing essential features. In this context, we focus exclusively on terms up to the $(1-y)^{2}$ order, disregarding higher-order terms in Eq. (160). This approach is sufficient to explain the interesting results discussed in the later section IV. To accomplish this, we again approximate the quantity $\Bigl{(}1+\frac{P(p_{\parallel},y)}{\omega(\bm{p},y)}\Bigr{)}$ using the Taylor series expansion near $y\rightarrow 1$. This yields:
@@
Therefore, we can express the particle distribution function approximately in terms of the small parameter $(1-y)$ up to the second order. The expression is given by:
@@
$\mathrm{C}_{1}(\bm{p},y)$ $=4y\Omega_{0}|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau(P_{1}-\omega_{1})\cos{(\Upsilon)}+4y|\Gamma_{2}|^{2}(2\Omega_{0}(\omega_{1}^{2}-\omega_{0}\omega_{1}+(\omega_{1}-P_{1})E_{0}\tau)+\Omega_{1}\omega_{1}^{2}y),$ (167)
@@
$+16E_{0}^{2}\tau^{2}\epsilon^{2}_{\perp}(1+\tau^{2}\omega_{1}^{2})\biggr{)}\Bigr{]}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|\cos{(\Upsilon)}}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}\Bigl{(}-16E_{0}^{4}\Omega_{0}\omega_{1}^{4}\tau^{8}y^{2}$
@@
$-4\Omega_{0}\omega_{1}(4+(-\omega_{0}^{2}+\omega_{1}^{2})\tau^{2})y^{2})\Bigr{)}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|\Omega_{0}\tau y}{(\omega_{1}(4+\omega_{1}^{2}\tau^{2})))}\Bigl{(}2(4+\omega_{1}^{2}\tau^{2})(\omega_{1}^{3}(2+\omega_{1})$
@@
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Bigr{]}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2}).$ (173)
Figure 20: Total Probability of pair production at time $t=15[m^{-1}],$ where $-3<\nu<3,E_{0}=0.2E_{c},$ and $\tau=10[m^{-1}].$
A.2 Fourier analysis

### v6

@@ VI Conclusion
We have conducted a detailed analysis of the electron-positron pair production from the vacuum under the influence of a time-dependent Sauter-pulse electric field. We calculated the particle distribution function using the exact analytic solution for the mode function. We have developed an analytical theory that is valid for finite times $t>\tau.$ In particular, we find an analytical expression for the distribution function in the power series of the small parameter (1y).(1-y). The interesting dynamical features of the momentum distribution function at finite times are attributed to the function that appears in this expansion. We have further elaborated on this aspect of the finite-time behavior of the longitudinal and transverse momentum spectra.
@@
We analyzed the temporal evolution of the quasiparticle distribution function. Our investigation revealed that the process of transition from initially virtual particles to real particles occurs in three distinct stages, which crucially depend on the longitudinal and transverse momentum components. Moreover, we have quantified the initiation of the REPP stage, determined by the momentum value, with higher momenta resulting in narrower oscillations in the transient region. We meticulously examined the LMS and TMS to understand the momentum-dependent behavior during these stages. In the LMS, one interesting feature at the beginning of the REPP stage is that the spectrum exhibits an oscillating structure as imprints of a quantum signature at the finite time when the electric field is nearly zero. The two-peaked structure, with the central Gaussian-like peak exhibiting onset oscillations, evolves and eventually fades away. Based on this observation, we identified three distinct time scales associated with this behavior. Additionally, we investigated the impact of different electric field strengths ($E_{0}$) on the duration of the interference pattern’s formation and disappearance. In the multiphoton regime, for $\gamma=2.5$, we explored the time evolution of LMS in different stages of pair production, and the spectrum shows the splitting of a smooth uni-modal structure into a multi-modal Gaussian structure near the REPP region, after which it merges into a single peak Gaussian profile at the asymptotic time limit.
@@
Utilizing the approximate expressions for the momentum distribution function, we have shown that the LMS structure arises from three distinct functional behaviors. The second-order term dominantly governs the early times, while combining the first and zeroth-order terms leads to the central peak structure with onset oscillation. The first-order term contributes to oscillations, owing to the presence of a Gaussian envelope with oscillatory behavior, but its amplitude diminishes with time, resulting in a smooth spectrum profile at late times.
@@
We examined the role of vacuum polarization $u(p_{\parallel},t)$, and its counterpart $v(p_{\parallel},t)$ by plotting their time evolution. In the REPP region, both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ exhibit nearly identical oscillations with the same amplitude, and this amplitude of oscillation decays for higher $p_{\parallel}$ values. This observation implies that the qualitative nature of the vacuum polarization function is significantly influenced by the longitudinal momentum. To elucidate this further, we plotted the LMS of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ These plots reveal two distinct patterns emerging at the start of the REPP stage. One pattern exhibits a Gaussian-like structure, while the other displays a deformed oscillatory profile within the Gaussian envelope. These oscillations, exhibiting varying amplitudes, are confined to a small window of longitudinal momentum ($1<p_{\parallel}<1$).Over time, the regular oscillations of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ became balanced. This balance is explicitly observed in the late-time LMS of the created particles, where oscillations are absent due to the equilibrium between $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ As we recognize the pivotal roles of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ functions in accelerating (creation of $e^{-}$) and decelerating (annihilation of $e^{-}$) electrons, respectively, the observed oscillations in the $f(p_{\parallel},t)$ at finite times can be attributed to the characteristic behavior imprints of these functions, resulting from the acceleration and deceleration of electrons in momentum representation, which manifests as the observed oscillating structure. Next, we also discussed the influence of transverse momentum on LMS, which diminishes the value of $f(p_{\parallel})$ and smooths out the oscillation for higher $p_{\perp}$ values. Moreover, we emphasized that as transverse momentum increases, these oscillations in the late REPP stage are smoothed, and the pair formation time is affected.
@@
Finally, we studied the dynamics of TMS, uncovering interesting results that have received limited attention from other authors, particularly in the context of the Sauter pulsed electric field. We observe fluctuating substructures that undergo regular changes in the QEPP and transient region depending on the momentum value. This substructure evolution is further influenced by longitudinal momentum, leading to distinct spectral shapes at given times. This finding indicates that the longitudinal and transverse momentum are related to each other during particle-antiparticle formation quantitatively and qualitatively. At the beginning of the REPP region, this substructure disappears. The shape of the spectrum becomes an invariable Gaussian-like profile and is not affected by the longitudinal momentum, but its peak value still depends on longitudinal momentum during the REPP stage. This contrast highlights the dynamic evolution of substructures in earlier stages, in contrast to the smooth Gaussian-like profile observed in the REPP stage. A similar intriguing behavior in the transverse momentum distribution has been observed in the works by Krajewska et al. Krajewska and Kamiński (2019), Bechler et al. Bechler et al. (2023), and Majczak et al. Majczak et al. (2024), depending on the specifics of the electric field parameters. These studies highlight a fascinating characteristic of the transverse distribution function peak that occurs at non-zero $p_{\perp}$ value depending on the frequency. This observation prompts questions about the feasibility of pair formation through tunneling theory, particularly considering the significant distribution function values at non-zero transverse momentum for low frequencies. In our case, the spectrum also shows a peak at non-zero transverse momentum, which signifies pair creation through multiphoton processes at finite times.
The finite-time analysis of LMS and TMS reveals not only the critical momentum dependence of the pair creation dynamics but also provides direct insight into observable phenomena. Crucially, as established in Sec.V, the quasiparticle distribution at time $T$ corresponds to measurable pair production if the field is switched off at that instant. Thus, the oscillatory structures and momentum spectra analyzed here represent genuine physical effects.
Moreover, investigating the LMS and TMS of the created quasiparticles at different dynamical stages demonstrates a pronounced dependence on both longitudinal and transverse momenta, while simultaneously uncovering the nontrivial behavior of the vacuum during finite-time pair production processes.
VII Acknowledgments
@@
We are grateful to the anonymous referee for constructive comments that helped improve the manuscript. Deepak gratefully acknowledge the financial support from Homi Bhabha National Institute (HBNI) for carrying out this research work.
Appendix A Evaluation approximate analytical expression for the distribution function
@@
[…]
@@
$|\Gamma(\textrm{i}z)|^{2}=\frac{\pi}{z\sinh{(\pi z)}},$ $|\Gamma(1+\textrm{i}z)|^{2}=\frac{\pi z}{\sinh{(\pi z)}},$ $|\Gamma(\frac{1}{2}+\textrm{i}z)|^{2}$ $=\frac{\pi}{\cosh{(\pi z)}}.$ (167)
@@
When computing expressions like $\Gamma_{1}\Gamma^{*}_{2}$, approximate methods prove advantageous. A frequently utilized technique entails utilizing Stirling’s formula for the Gamma function, offering a simpler yet effective approach to assess the desired expression.
@@
$\Gamma_{1}\Gamma^{*}_{2}$ $\approx|\Gamma_{1}\Gamma^{*}_{2}|e^{\textrm{i}\mathrm{\varrho}}$ (171)
@@
$|\Gamma_{1}\Gamma^{*}_{2}|$ $=\frac{\omega_{0}}{2\omega_{1}}\frac{\exp{(\pi\tau/2(\omega_{0}-\omega_{1}+2E_{0}\tau))}}{\sinh{(\pi\tau\omega_{0})}\sqrt{1+\omega_{1}^{2}\tau^{2}}}\sqrt{\frac{((\omega_{0}+2E_{0}\tau)^{2}-\omega_{1}^{2})}{(\omega_{1}^{2}-(2E_{0}\tau-\omega_{0})^{2})}},$ (172)
@@
We aim to express $f(\bm{p},y)$ in a series involving (1y)(1-y). We can then consider truncating higher-order terms to simplify the analysis while still capturing essential features. In this context, we focus exclusively on terms up to the $(1-y)^{2}$ order, disregarding higher-order terms in Eq. (174). This approach is sufficient to explain the interesting results discussed in the later section IV. To accomplish this, we again approximate the quantity $\Bigl(1+\frac{P(p_{\parallel},y)}{\omega(\bm{p},y)}\Bigr)$ using the Taylor series expansion near $y\rightarrow 1$. This yields:
@@
Therefore, we can express the particle distribution function approximately in terms of the small parameter (1y)(1-y) up to the second order. The expression is given by:
@@
$\mathrm{C}_{1}(\bm{p},y)$ $=4y\Omega_{0}|\Gamma_{1}\Gamma^{*}_{2}|E_{0}\tau(P_{1}-\omega_{1})\cos{(\Upsilon)}+4y|\Gamma_{2}|^{2}(2\Omega_{0}(\omega_{1}^{2}-\omega_{0}\omega_{1}+(\omega_{1}-P_{1})E_{0}\tau)+\Omega_{1}\omega_{1}^{2}y),$ (181)
@@
$+16E_{0}^{2}\tau^{2}\epsilon^{2}_{\perp}(1+\tau^{2}\omega_{1}^{2})\biggr)\Bigr]+\frac{|\Gamma_{1}\Gamma^{*}_{2}|\cos{(\Upsilon)}}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}\Bigl(-16E_{0}^{4}\Omega_{0}\omega_{1}^{4}\tau^{8}y^{2}$
@@
$-4\Omega_{0}\omega_{1}(4+(-\omega_{0}^{2}+\omega_{1}^{2})\tau^{2})y^{2})\Bigr)+\frac{|\Gamma_{1}\Gamma^{*}_{2}|\Omega_{0}\tau y}{(\omega_{1}(4+\omega_{1}^{2}\tau^{2})))}\Bigl(2(4+\omega_{1}^{2}\tau^{2})(\omega_{1}^{3}(2+\omega_{1})$
@@
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Bigr]+\frac{|\Gamma_{1}\Gamma^{*}_{2}|}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2}).$ (187)
Figure 20: Total Probability of pair production at time $t=15[m^{-1}],$ where $3<\nu<3,E_{0}=0.2E_{c},$ and $\tau=10[m^{-1}].$
Appendix B Fourier analysis
@@
Appendix C Explicit Spinor solutions for free and Sauter-pulse fields
This appendix provides the explicit forms of the Dirac spinor solutions used in Section V for the free-field regions (I and III) and the Sauter-pulse region (II), as well as the details of the matching procedure at the switch-off time $t=T.$
C.0.1 Free-field spinors (Regions I and III)
In regions where the electric field vanishes, the Dirac equation admits plane-wave solutions of the form Avetissian (2016)
$\Psi_{\bm{p},r}^{(\lambda)}(\bm{x},t)=\xi_{\bm{p},r}^{(\lambda)}(t)e^{i\bm{p}\cdot\bm{x}},\quad\lambda=\pm,$
where $\lambda=+$ corresponds to positive-energy (electron) states and $\lambda=$- to negative-energy (positron) states. The canonical momentum $\bm{p}=(p_{1},p_{2},p_{\parallel})$ is conserved, and the kinetic momentum along the field direction is $\Pi_{3}=p_{\parallel}-eA$ with $A$ constant. The corresponding energy is
$\mathcal{E}=\sqrt{m^{2}+p_{\perp}^{2}+\Pi_{3}^{2}},\qquad p_{\perp}^{2}=p_{1}^{2}+p_{2}^{2}.$
For spin projection $r=1$, the normalized free spinors are
$\xi_{\bm{p},1}^{(+)}(t)$ $=\frac{1}{\sqrt{2\mathcal{E}(\mathcal{E}-\Pi_{3})}}\begin{pmatrix}m\\[2.0pt] 0\\[2.0pt] -\mathcal{E}+\Pi_{3}\\[2.0pt] p_{1}+ip_{2}\end{pmatrix}e^{-i\mathcal{E}t},$ (191)
$\xi_{\bm{p},1}^{(-)}(t)$ $=\frac{1}{\sqrt{2\mathcal{E}(\mathcal{E}+\Pi_{3})}}\begin{pmatrix}m\\[2.0pt] 0\\[2.0pt] \mathcal{E}+\Pi_{3}\\[2.0pt] p_{1}+ip_{2}\end{pmatrix}e^{i\mathcal{E}t}.$ (192)
For $r=2$, the corresponding spinors are
$\xi_{\bm{p},2}^{(+)}(t)$ $=\frac{1}{\sqrt{2\mathcal{E}(\mathcal{E}+\Pi_{3})}}\begin{pmatrix}p_{1}-ip_{2}\\[2.0pt] \mathcal{E}+\Pi_{3}\\[2.0pt] 0\\[2.0pt] m\end{pmatrix}e^{-i\mathcal{E}t},$ (193)
$\xi_{\bm{p},2}^{(-)}(t)$ $=\frac{1}{\sqrt{2\mathcal{E}(\mathcal{E}-\Pi_{3})}}\begin{pmatrix}-p_{1}+ip_{2}\\[2.0pt] \mathcal{E}-\Pi_{3}\\[2.0pt] 0\\[2.0pt] m\end{pmatrix}e^{i\mathcal{E}t}.$ (194)
C.0.2 Sauter-pulse spinors (Region II)
In Region II, the electric field is given by the Sauter pulse $E(t)=E_{0}\operatorname{sech}^{2}(t/\tau)$. The Dirac spinor is constructed from the scalar mode function $\psi_{\bm{p}}^{(\pm)}(t)$ obtained in Section IV. Using the constant spinor basis vectors $R_{r}$ defined in Eq. (12), the positive-energy solution is
$\Psi_{\bm{p},r}^{(+)}(t)=\big[i\gamma^{0}\partial_{t}+\bm{\gamma}\cdot\bm{p}-e\gamma^{3}A(t)+m\big]\psi_{\bm{p}}^{(+)}(t)R_{r}\,.$ (195)
For $r=1$, this takes the explicit form
$\Psi_{\bm{p},1}^{(+)}(t)=\begin{pmatrix}m\\ 0\\ i\partial_{t}+P(p_{\parallel},t)\\ p_{1}+ip_{2}\end{pmatrix}\psi_{\bm{p}}^{(+)}(t),\qquad P(p_{\parallel},t)=p_{\parallel}-eA(t),$ (196)
with the mode function $\psi_{\bm{p}}^{(+)}(t)$ given by Gauss-hypergeometric expression. The corresponding negative-energy spinor $\Psi_{\bm{p},1}^{(-)}(t)$ is obtained by replacing $\psi_{\bm{p}}^{(+)}(t)$ with $\psi_{\bm{p}}^{(-)}(t)$ and changing the sign of the energy component.
C.0.3 Matching at $t=T$ and the reflection amplitude
In Section V, the field is switched off abruptly at $t=T$. The solution in Region II must match continuously onto a superposition of free states in Region III:
$\Psi_{\bm{p},1}^{(+)}(T)=\mathcal{R}(\bm{p})\,\xi_{\bm{p},1}^{(-)}(T)+\mathcal{I}(\bm{p})\,\xi_{\bm{p},1}^{(+)}(T),$ (197)
where $\mathcal{R}(\bm{p})$ is the pair-production amplitude (reflection from the Dirac sea) and $\mathcal{I}(\bm{p})$ is the transmission amplitude. Projecting both sides onto $\xi_{\bm{p},1}^{(-)}(T)$ using orthonormality yields
$\mathcal{R}(\bm{p})=\big[\xi_{\bm{p},1}^{(-)}(T)\big]^{\dagger}\Psi_{\bm{p},1}^{(+)}(T).$
Substituting the explicit forms of $\xi_{\bm{p},1}^{(-)}$ and $\Psi_{\bm{p},1}^{(+)}$ leads to Eq. (147) in the main text, which after simplification gives the probability $\mathcal{W}(\bm{p},T)=2|\mathcal{R}(\bm{p})|^{2}$.
The identity $f(\bm{p},T)=\mathcal{W}(\bm{p},T)$ established in Section V follows from comparing this expression with the quasiparticle distribution function $f(\bm{p},t)$ derived in Section III.1.